The temperature coefficient of resistance for carbon at 0oC is -0.0048/oC. What is the significance of the minus sign? A carbon resistor has a resistance of 500Ω.Determine its resistance at 50oC.
1
Expert's answer
2020-05-04T12:43:56-0400

"R_t=R_0(1+\alpha \triangle t)"

"R_t" is the resistance at a particular temperature.

Minus sign of temperature coefficient "(\alpha)" means that the resistance will decrease with increase in temperature.

Resistance increases, in case of conductor, with increase in temperature and decreases, in case of insulator, due to increase in temperature. 

It shows that carbon is an insulator.

"R_{50}=R_0(1+\alpha \times 50)"

"\implies R_{50}=500(1-0.0048\times 50)=380\Omega"